Table 2.1 contains information on three techniques for producing $15 worth of bar soap. Assume that we said $15 worth of bar soap" because soap costs $3 per bar and all three techniques produce 5 bars of soap ($15 $3 per bar x 5 bars). So you know each technique produces 5 bars of soap. LO2.3 a. What technique will you want to use if the price of a bar of3 soap falls to $2.75? What if the price of a bar of soap rises to $4? To $5? 


Where do I substitute the $2.75


1
Expert's answer
2021-08-20T18:04:04-0400

production cost will be

"3\space \times 2.75=\$8.25"

"3\space \times 4=\$12"

"3\space \times 5=\$15"

I will go for technique with $2.75 because it has low production cost.